# Dynamic angular synchronization under smoothness constraints

### v1

@@ 1.2 Dynamic angular synchronization
      Global matrix denoising and local synchronization (GMD-LTRS-DynSync; Algorithm 3): Projects the measurement matrix onto the low-frequency eigenspace of a smoothness operator (matrix denoising), followed by locally solving (at each time $k$) the spectral relaxation of the angular synchronization problem.
  • We provide theoretical guarantees for Algorithm 3, in the form of non-asymptotic high probability bounds for the global $\ell_{2}$ estimation error, under the dynamic versions of the AGN and Outliers model (see Corollaries 1 and 2 respectively). In particular, under mild smoothness assumptions, we prove that (1.3) holds for the output of Algorithm 3, and give precise non-asymptotic bounds from which the convergence rates can be deduced.
@@
      In the case of the Outliers model, our results apply even when the measurement graphs are very sparse – each being potentially disconnected (see Remarks 8 and 10).
@@
      For the AGN model, our results show that Algorithm 3 succeeds even when the noise level $\sigma$ is large w.r.t $n$, and is allowed to grow with $T$ as $O(T^{\alpha})$ for $\alpha\in[0,1/4)$ (see Remarks 5 and 9).
  • Finally, we support our theoretical findings with extensive experiments on synthetic data, for the aforementioned AGN and Outliers models. Notably, our experiments demonstrate that Algorithm 3, which is the focus of our theoretical analysis, usually outperforms other competing methods in terms of accuracy.
@@ 1.3 Related work
  • The work [3] considered the dynamic translation synchronization problem which is related to the BTL model for ranking. Here, the latent strength vector $r^{*}(k)\in\mathbb{R}^{n}$ ($k=1,2,\dots,T$) satisfies a similar smoothness assumption as (1.2). Given a measurement graph $\mathcal{H}_{k}=([n],\mathcal{E}_{k})$ at time $k$, we are given noisy information $r_{i}^{*}(k)-r^{*}_{j}(k)+\epsilon_{ij}(k)$ for each $\left\{{i,j}\right\}\in\mathcal{E}_{k}$, where $\epsilon_{ij}(k)$ is zero-mean subgaussian noise. Two estimators were proposed, the first of which involves solving an unconstrained smoothness-penalized least squares problem. The second estimator involves first estimating $r^{*}(k)$ via least-squares minimization, at each $k$, and then projecting the vector of estimated solutions on to the low-frequency space of a smoothness operator. High-probability bounds were obtained for the MSE, establishing their consistency when $T$ is large and the rest of parameters are fixed. In particular, these results imply that the proposed estimators can tolerate larger levels of noise compared to the static case. Our Algorithms 1 and 2 are motivated from these two methods, the key difference being that we now need to incorporate the (relaxations of) the constraints, imposed by the group $\text{SO}(2)$. This in particular leads to solving a TRS which is challenging to analyze theoretically. We also remark that the analysis in [3] requires each measurement graph to be connected due to technical reasons. Our analysis for Algorithm 3, however, does not suffer from these issues, and applies even when the graphs are very sparse and disconnected. We believe that a similar procedure can be applied for the dynamic translation synchronization problem, leading to similar guarantees as in the present paper (Theorem 2 and Corollary 2).
@@ 1.4 Paper structure
The remainder of this paper is structured as follows. In Section 2 we present the setup for dynamic synchronization. In Section 2.2 we introduce the dynamic versions of the AGN and Outliers model, togheter with our global smoothness assumptions. Our proposed estimators and the algorithms to compute them are described in Section 2.3. Section 3, which consists on the theoretical analysis of Algorithm 3, is divided in three parts. In Sections 3.1 and 3.2 we analyse the matrix denoising stage for the AGN and Outliers model respectively. In Section 3.3 we provide an analysis for the local synchronization stage of Algorithm 3, which valid for both noise models. Section 4 contains numerical experiments on synthetic data and Section 5 contains concluding remarks.
@@ 2.2 Problem setup
Denoting $M\in\mathbb{R}^{T\times(T-1)}$ to be the edge-incidence matrix of the path graph on $T$ vertices, this can be written equivalently as
${g^{*}}^{\mathsf{H}}((MM^{\top})\otimes I_{n})g^{*}\leq S_{T}.$
The above assumption states that $g^{*}(k)$ does not change too quickly with $k$ on “average”, and resembles the notion of quadratic variation of a function. Since Assumption 1 is in place, we do not need to worry about the invariance of (2.4) with respect to constant shifts.
@@ 2.3.3 Denoising measurement matrices
  • Local synchronization stage. After “Hermitianizing” each block of $\widehat{G}$, we solve, for each $k\in[T]$, a “local” synchronization problem, as in (2.14), but with input data $\widehat{G}(k)$. In Algorithm 3, we solve its TRS relaxation as in (2.16).
@@
    Find $\widehat{g}^{\prime}(k)\leftarrow$ the eigenvector associated to the largest eigenvalue of $\widehat{G}(k)$.
@@
    Define $\widehat{g}(k)=\mathcal{P}_{\mathcal{C}_{n}}(\widehat{g}^{\prime}(k))e^{-\iota\hat{\theta}_{1}(k)}$, where $\hat{\theta}_{1}(k)$ is defined by $\widehat{g}^{\prime}_{1}(k)=|\widehat{g}^{\prime}_{1}(k)|e^{\iota\hat{\theta}_{1}(k)}$.
While the spectral method may sacrifice some accuracy compared to the TRS-based approach, it offers simplicity in analysis. Our theoretical analysis in Section 3 is based on the spectral approach, whereas in our experiments, we opt for the TRS-based method.
@@
4:     Hermitianize $\widehat{G}(k)\leftarrow(\widehat{G}(k)+{\widehat{G}(k)}^{\mathsf{H}})/2$.
5:     Find $\widehat{\tilde{g}}(k)\leftarrow$ solution of (2.16) with data $\widehat{G}(k)$.
@@ 3 Analysis
In this section, we present theoretical guarantees for Algorithm 3 under the AGN and Outliers models. Sections 3.1 and 3.2 contain results specifically addressing the matrix denoising stage of Algorithm 3 (as elaborated in Section 2.3.3). Then, Section 3.3 uses these estimates to provide guarantees for recovering $g^{*}(1),\dots,g^{*}(T)$ – it provides an analysis of the local synchronization stage by employing the spectral approach outlined in Remark 3.
@@ 3.1 Matrix denoising under AGN model
Theorem 1.
@@ 3.2 Matrix denoising under the Outliers model
Theorem 2.
@@ 3.3 Spectral local synchronization guarantee
We now proceed to establish guarantees for the local synchronization phase within Algorithm 3. While Algorithm 3 employs a TRS-based methodology at this stage, it is easier to analyze the spectral method, as delineated in Remark 3. Our strategy involves utilizing the Davis-Kahan inequality “locally”, focusing on specific time instances where the error converges to zero.
To begin with, denote for each $k$ a scaling $s(k)\geq 0$ where $s(k)=1$ for the AGN model, and $s(k)=d(k)=(1-\eta)p(k)$ for the outliers model (recall Assumption 3). Then note that since $\overline{G}^{*}(k)$ is symmetric, if $\widehat{G}^{\prime}(k):=\frac{1}{2}(\widehat{G}(k)+{\widehat{G}}^{\mathsf{H}}(k))$, we have
@@
The next step is to use the Davis-Kahan theorem [19] to relate the largest eigenvector of $\widehat{G}^{\prime}(k)$ to $g^{*}(k)/\sqrt{n}$, for each $k$. Since $s(k)\overline{G}^{*}(k)=s(k)(g^{*}(k){g^{*}(k)}^{\mathsf{H}}-I_{n})$, the spectral gap between the largest and second largest eigenvalues of $s(k)\overline{G}^{*}(k)$ is $s(k)n$. Following the notation in Section 2.3.3, denote $\widehat{g}^{\prime}(k)$ to be the eigenvector of $\widehat{G}^{\prime}(k)$ associated with its largest eigenvalue. Then, by the Davis-Kahan theorem [19], we have for each $k$ that
@@
Let $g^{*}$ be a ground truth vector satisfying Assumption 1 as defined in (2.1). Denote $\widehat{g}^{\prime}(k)$ to be the largest eigenvector of $\widehat{G}^{\prime}(k)$ and $\widehat{g}(k):=\mathcal{P}_{\mathcal{C}_{n}}(\widehat{g}^{\prime}(k))e^{-\iota\hat{\theta}_{1}(k)}$, where $\hat{\theta}_{1}(k)$ is defined by $\widehat{g}^{\prime}_{1}(k)=|\widehat{g}^{\prime}_{1}(k)|e^{\iota\hat{\theta}_{1}(k)}$ (with $\hat{\theta}_{1}(k)=0$ if $\widehat{g}^{\prime}_{1}(k)=0$). Let $\widehat{g}=\operatorname{concat}\left(\widehat{g}(1),\ldots,\widehat{g}(T)\right)$. Then there exists a constant $C>0$ such that
@@
Corollary 1.
Let $A$ be the stacked measurement matrix for the AGN model in (2.2), with a ground truth $g^{*}$ satisfying Assumption 2. Choosing $\tau=\tau^{*}$ with $\tau^{*}$ as in (3.4), consider the estimator $\widehat{g}=\operatorname{concat}\left(\widehat{g}(1),\ldots,\widehat{g}(T)\right)$ defined in Proposition 1. There exist constants $C_{1},C_{2},C_{3},C_{4}>0$ such that for any $\delta\in(0,1)$ the following is true.
@@
    In particular, when $\sigma>0$ is a constant, then $\|\widehat{g}-g^{*}\|_{2}^{2}=O(T^{2/3}S_{T}^{1/3})$ which matches the rate for denoising smooth signals on a path graph [40, Theorem 6].
Corollary 2.
Let $A$ be the stacked measurement matrix under the outliers model (2.3), with a ground truth $g^{*}$ satisfying Assumption 2, and the matrix $D$ (c.f. (3.9)) satisfies Assumption 3 with parameter $\mu\geq 1$. For $\tau=\tau^{*}$ as in (3.12), consider the estimator $\widehat{g}=\operatorname{concat}\left(\widehat{g}(1),\ldots,\widehat{g}(T)\right)$ defined in Proposition 1 and recall the notation of Theorem 2. There exist constants $C_{1}\in(0,1)$ and $C_{2},C_{3},C_{4}>0$ such that for any $\delta\in(0,C_{1})$, the following is true.
@@
    Consider the setup of Remark 8 so that $p_{\min}\asymp 1/n$. Then if $T$ satisfies (3.16) (which is the case for $T$ large enough when $S_{T}=\omega(1/T^{2})$ and $S_{T}=o(T)$ holds), then we obtain the bound

### v2

@@ 1.2 Dynamic angular synchronization
      Global matrix denoising and local synchronization (GMD-LTRS-DynSync; Algorithm 3): Projects the measurement matrix onto the low-frequency eigenspace of a smoothness operator (matrix denoising), followed by locally solving (at each time $k$) the sphere-relaxation of the angular synchronization problem via a TRS formulation.
  • As the TRS estimator is hard to analyze theoretically, we analyze a slight modification of Algorithm 3, namely Algorithm 4, where the TRS estimator for the local synchronization step is replaced by an alternative spectral approach (see also Remark 3). We provide theoretical guarantees for Algorithm 4, in the form of non-asymptotic high probability bounds for the global $\ell_{2}$ estimation error, under the dynamic versions of the AGN and Outliers model (see Corollaries 1 and 2 respectively). In particular, under mild smoothness assumptions, we prove that (1.3) holds for the output of Algorithm 4, and give precise non-asymptotic bounds from which the convergence rates can be deduced.
@@
      In the case of the Outliers model, our results apply even when the measurement graphs are very sparse – each being potentially disconnected (see Remarks 10 and 12).
@@
      For the AGN model, our results show that Algorithm 4 succeeds even when the noise level $\sigma$ is large w.r.t $n$, and is allowed to grow with $T$ as $O(T^{\alpha})$ for $\alpha\in[0,1/4)$ (see Remarks 7 and 11).
  • Finally, we support our theoretical findings with extensive experiments on synthetic data, for the aforementioned AGN and Outliers models. Notably, our experiments demonstrate that Algorithm 3 usually outperforms other competing methods in terms of accuracy.
@@ 1.3 Related work
  • The work [3] considered the dynamic translation synchronization problem which is related to the BTL model for ranking. Here, the latent strength vector $r^{*}(k)\in\mathbb{R}^{n}$ ($k=1,2,\dots,T$) satisfies a similar smoothness assumption as (1.2). Given a measurement graph $\mathcal{H}_{k}=([n],\mathcal{E}_{k})$ at time $k$, we are given noisy information $r_{i}^{*}(k)-r^{*}_{j}(k)+\epsilon_{ij}(k)$ for each $\left\{{i,j}\right\}\in\mathcal{E}_{k}$, where $\epsilon_{ij}(k)$ is zero-mean subgaussian noise. Two estimators were proposed, the first of which involves solving an unconstrained smoothness-penalized least squares problem. The second estimator involves first estimating $r^{*}(k)$ via least-squares minimization, at each $k$, and then projecting the vector of estimated solutions on to the low-frequency space of a smoothness operator. High-probability bounds were obtained for the MSE, establishing their consistency when $T$ is large and the rest of parameters are fixed. In particular, these results imply that the proposed estimators can tolerate larger levels of noise compared to the static case. Our Algorithms 1 and 2 are motivated from these two methods, the key difference being that we now need to incorporate the (relaxations of) the constraints, imposed by the group $\text{SO}(2)$. This in particular leads to solving a TRS-based sphere-relaxation which is challenging to analyze theoretically. We also remark that the analysis in [3] requires each measurement graph to be connected due to technical reasons. Our analysis for Algorithm 4, however, does not suffer from these issues, and applies even when the graphs are very sparse and disconnected. We believe that a similar procedure can be applied for the dynamic translation synchronization problem, leading to similar guarantees as in the present paper (Theorem 2 and Corollary 2).
@@ 1.4 Paper structure
The remainder of this paper is structured as follows. In Section 2 we present the setup for dynamic synchronization. In Section 2.2 we introduce the dynamic versions of the AGN and Outliers model, togheter with our global smoothness assumptions. Our proposed estimators and the algorithms to compute them are described in Section 2.3. Section 3, which consists of the theoretical analysis of Algorithm 4, is divided into three parts. In Sections 3.1 and 3.2 we analyse the matrix denoising stage for the AGN and Outliers model respectively. In Section 3.3 we provide an analysis for the local synchronization stage of Algorithm 4, which is valid for both noise models. Section 4 contains numerical experiments on synthetic data and Section 5 contains concluding remarks.
@@ 2.2 Problem setup
Denoting $M\in\left\{{0,1,-1}\right\}^{T\times(T-1)}$ to be the incidence matrix of any directed path graph [^2] The rows of $M$ correspond to the $T$ vertices, and the columns correspond to the $T-1$ directed edges. For the column corresponding to edge $i\rightarrow j$, we set its $i$th entry to $+1$, the $j$th entry to $-1$, and the rest of the entries in that column to 00. on $T$ vertices, this can be written equivalently as
${\color[rgb]{0,0,0}\|{(M^{\top}\otimes I_{n})g^{*}}\|_{2}^{2}}={g^{*}}^{\mathsf{H}}((MM^{\top})\otimes I_{n})g^{*}\leq S_{T},$
where $MM^{\top}$ corresponds to the Laplacian matrix of the undirected path graph.
The above assumption states that $g^{*}(k)$ does not change too quickly with $k$ on “average”, and resembles the notion of quadratic variation of a function. It can also be viewed as stating that $g^{*}$ lies “close” (depending on $S_{T}$) to the subspace spanned by the “smallest few” eigenvectors of $(MM^{\top})\otimes I_{n}$. The closed-form expressions of the eigenvectors (and also eigenvalues) of $MM^{\top}$ are well-known (e.g., [10]), and can be viewed as the Fourier basis for signals residing on that graph (with the smallest eigenvector corresponding to the lowest frequency signal, and so on). Since Assumption 1 is in place, we do not need to worry about the invariance of (2.4) with respect to constant shifts.
@@ 2.3.3 Denoising measurement matrices
  • Local synchronization stage. For each $k\in[T]$ we first “Hermitianize” the block $\widehat{G}(k)$ to obtain $\widehat{G}^{\prime}(k)$, and then solve a “local” synchronization problem, as in (2.14), but with input data $\widehat{G}^{\prime}(k)$. In Algorithm 3, we solve its TRS relaxation as in (2.16).
@@
    Find $\widehat{g}^{\prime}(k)\leftarrow$ the eigenvector associated to the largest eigenvalue of $\widehat{G}^{\prime}(k)$.
@@
    Define $\widehat{g}(k)=\mathcal{P}_{\mathcal{C}_{n}}(\widehat{g}^{\prime}(k))e^{-\iota\hat{\theta}_{1}(k)}$, where $\hat{\theta}_{1}(k)$ is defined by $\widehat{g}^{\prime}_{1}(k)=|\widehat{g}^{\prime}_{1}(k)|e^{\iota\hat{\theta}_{1}(k)}$ (with $\hat{\theta}_{1}(k)=0$ if $\widehat{g}^{\prime}_{1}(k)=0$).
While the spectral method may sacrifice some accuracy compared to the TRS-based approach, it offers simplicity in analysis. Our theoretical analysis in Section 3 is based on the spectral approach (outlined separately as Algorithm 4 for clarity), whereas in our experiments, we opt for the TRS-based method.
Remark 4 (Denoising smooth signals on a graph).
As alluded to earlier, the matrix denoising step in Algorithm 3 is motivated by the Laplacian eigenmaps estimator (see e.g., [40]) for denoising smooth signals on a graph. To make the connection explicit using our notation, we recall that in this latter problem, we are given an undirected graph $\mathcal{H}=([T],\mathcal{E})$, along with noisy measurements $y=x^{*}+\eta$ of a signal $x^{*}\in\mathbb{R}^{T}$ with $\eta$ denoting the zero-mean noise. Denoting $L$ to be the Laplacian matrix of $\mathcal{H}$, here $x^{*}$ is assumed to be smooth w.r.t $\mathcal{H}$ in the sense that
$(x^{*})^{\top}Lx^{*}=\sum_{\left\{{i,j}\right\}\in\mathcal{E}}(x^{*}_{i}-x^{*}_{j})^{2}\leq S_{T}.$ (2.17)
The Laplacian eigenmaps estimator then estimates $x^{*}$ by projecting $y$ on to the space spanned by the $\tau$ smallest eigenvectors of $L$.
In our setting, the aforementioned graph $\mathcal{H}$ is a path graph333since we consider its vertices to denote time instants, although the analysis can be extended to general connected graphs as well. and each vertex $k\in[T]$ has a latent signal $g^{*}(k)\in\mathcal{C}_{n}$. We observe at each $k$ a noisy matrix $A(k)$ which is also in the form of a signal (scalar times $G^{*}(k)-I_{n}$) plus zero-mean noise term; recall (2.2) and (2.3). Moreover, $g^{*}(k)$’s satisfy a smoothness condition (recall (2.4)) which, as discussed earlier, translates to a smoothness condition on $G^{*}(k)$’s (recall (2.10)) that is analogous to (2.17).
As we will see in Section 3, the outline of the analysis is – at a top-level – similar to that in [40] in the sense of bounding the bias and variance error terms, and finding the optimal $\tau$. The difference lies in the technical details: we observe matrix-valued information (with complex-valued entries) at each node of $\mathcal{H}$ which leads to more cumbersome calculations – especially for the Outliers model – in deriving high probability error bounds.
Remark 5 (Comparison with [3]).
As discussed in Section 1.3, the work of [3] considered an estimator for the dynamic translation synchronization problem, which also relied on the idea of projecting on to the low-frequency subspace. In that work, the estimator first solved the least-squares problem locally at each time-point, and then projected the stacked vector of individual-solutions onto the low-frequency space. This required each individual graph to be connected for the analysis to work. In Algorithm 3, however, we first project the (stacked) input data matrices onto the low-frequency space, and then solve the local synchronization problem. This allows the individual graphs to be disconnected in the Outliers model – something we will prove in Section 3 for a modified version of Algorithm 3, namely Algorithm 4.
@@
4:     Hermitianize ${\color[rgb]{0,0,0}\widehat{G}^{\prime}(k):=}(\widehat{G}(k)+{\widehat{G}(k)}^{\mathsf{H}})/2$.
5:     Find $\widehat{\tilde{g}}(k)\leftarrow$ solution of (2.16) with data $\widehat{G}^{\prime}(k)$.
@@ 3 Analysis
In this section, we present theoretical guarantees for recovering $g^{*}(1),\dots,g^{*}(T)$ under the AGN and Outliers models. Sections 3.1 and 3.2 contain results specifically addressing the matrix denoising stage of Algorithm 3 (as elaborated in Section 2.3.3). The main results therein are Theorem 1 for the AGN model, and Theorem 2 for the Outliers model. Then, Section 3.3 uses these estimates to provide guarantees for recovering $g^{*}(1),\dots,g^{*}(T)$ – it provides an analysis of the local synchronization stage by employing the spectral approach outlined in Remark 3. The main results therein are Corollary 1 for the AGN model, and Corollary 2 for the Outliers model.
@@ 3.1 Matrix denoising under AGN model
Theorem 1 (Main result: $\overline{G}^{*}$ recovery under AGN model).
@@ 3.2 Matrix denoising under the Outliers model
Theorem 2 (Main result: $\overline{G}^{*}$ recovery under Outliers model).
@@ 3.3 Spectral local synchronization guarantee
We now proceed to establish guarantees for the local synchronization phase within Algorithm 3. While Algorithm 3 employs a TRS-based methodology at this stage, it is easier to analyze the spectral method, as delineated in Remark 3. This is outlined as a separate method in Algorithm 4 for clarity. Our main results for recovering $g^{*}$ are Corollaries 1 and 2 for the AGN and Outliers models respectively.
Algorithm 4 Global Matrix Denoising $+$ Local Spectral method for Dynamic Synchronization (GMD-LSPEC-DynSync)
3:for $k=1\dots,T$ do
4:     Hermitianize $\widehat{G}^{\prime}(k):=(\widehat{G}(k)+{\widehat{G}(k)}^{\mathsf{H}})/2$.
5:     Find $\widehat{g}^{\prime}(k)\leftarrow$ the eigenvector associated to the largest eigenvalue of $\widehat{G}^{\prime}(k)$.
6:     Find $\widehat{g}(k)=\mathcal{P}_{\mathcal{C}_{n}}(\widehat{g}^{\prime}(k))e^{-\iota\hat{\theta}_{1}(k)}$, where $\hat{\theta}_{1}(k)$ is defined by $\widehat{g}^{\prime}_{1}(k)=|\widehat{g}^{\prime}_{1}(k)|e^{\iota\hat{\theta}_{1}(k)}$ (with $\hat{\theta}_{1}(k)=0$ if $\widehat{g}^{\prime}_{1}(k)=0$).
7:end for
Our strategy involves utilizing the Davis-Kahan inequality “locally”, focusing on specific time instances where the error converges to zero. To begin with, denote for each $k$ a scaling $s(k)\geq 0$ where $s(k)=1$ for the AGN model, and $s(k)=d(k)=(1-\eta)p(k)$ for the outliers model (recall Assumption 3). Then note that since $\overline{G}^{*}(k)$ is symmetric, we have for the Hermitian matrix $\widehat{G}^{\prime}(k)$ defined in line $4$ of Algorithm 4 that
@@
The next step is to use the Davis-Kahan theorem [19] to relate the largest eigenvector of $\widehat{G}^{\prime}(k)$ to $g^{*}(k)/\sqrt{n}$, for each $k$. Since $s(k)\overline{G}^{*}(k)=s(k)(g^{*}(k){g^{*}(k)}^{\mathsf{H}}-I_{n})$, the spectral gap between the largest and second largest eigenvalues of $s(k)\overline{G}^{*}(k)$ is $s(k)n$. Recall $\widehat{g}^{\prime}(k)$ as defined in line $5$ of Algorithm 4. Then, by the Davis-Kahan theorem [19], we have for each $k$ that
@@
Let $g^{*}$ be a ground truth vector satisfying Assumption 1 as defined in (2.1), and let $\widehat{g}=\operatorname{concat}\big{(}\widehat{g}(1),\cdots,\widehat{g}(T)\big{)}\in\mathcal{C}_{nT}$ be obtained from Algorithm 4. Then there exists a constant $C>0$ such that
@@
Corollary 1 (Main result: $g^{*}$ recovery under AGN model).
Let $A$ be the stacked measurement matrix for the AGN model in (2.2), with a ground truth $g^{*}$ satisfying Assumption 2. Choosing $\tau=\tau^{*}$ with $\tau^{*}$ as in (3.4), consider the estimator $\widehat{g}=\operatorname{concat}\left(\widehat{g}(1),\ldots,\widehat{g}(T)\right)$ obtained in Algorithm 4. There exist constants $C_{1},C_{2},C_{3},C_{4}>0$ such that for any $\delta\in(0,1)$ the following is true.
@@
    In particular, when $\sigma>0$ is a constant, then $\|\widehat{g}-g^{*}\|_{2}^{2}=O(T^{2/3}S_{T}^{1/3})$ which matches the (optimal) rate for denoising smooth signals on a path graph on $T$ vertices [40, Theorems 5,6]; recall the discussion in Remark 4.
Corollary 2 (Main result: $g^{*}$ recovery under Outliers model).
Let $A$ be the stacked measurement matrix under the outliers model (2.3), with a ground truth $g^{*}$ satisfying Assumption 2, and the matrix $D$ (c.f. (3.9)) satisfies Assumption 3 with parameter $\mu\geq 1$. For $\tau=\tau^{*}$ as in (3.12), consider the estimator $\widehat{g}=\operatorname{concat}\left(\widehat{g}(1),\ldots,\widehat{g}(T)\right)$ obtained in Algorithm 4 and recall the notation of Theorem 2. There exist constants $C_{1}\in(0,1)$ and $C_{2},C_{3},C_{4}>0$ such that for any $\delta\in(0,C_{1})$, the following is true.
@@
    Consider the setup of Remark 10 so that $p_{\min}\asymp 1/n$. Then if $T$ satisfies (3.16) (which is the case for $T$ large enough when $S_{T}=\omega(1/T^{2})$ and $S_{T}=o(T)$ holds), we obtain the bound
@@
The bound exhibits the same scaling [^4] Note that the rates in [40, Theorems 5,6] are obtained for additive Gaussian noise. We believe the dependency on $T$ should be unchanged for more general sub-Gaussian noise models, however this needs to be established formally. w.r.t $T,S_{T}$ as in the AGN model; recall Remark 11.